OpenStudy (anonymous):

Please help! Will give medal and fan! If the length of each side of a cube is reduced to one-third its original value, is the surface area reduced to one-third its original value? Use an example to support your answer.

OpenStudy (anonymous):

no e.g. let lenght of each side=9 thensurface area=9*9=81 if length reduced by 1/3 then new lenght =3 then surface area=3*3=9 so its not 1/3 of 81 @keepsecure93